OpenStudy (nikolas):

Hi, I'm having trouble understanding a step in the solution for z^(n)+(1/(z^n))=2cos(nθ). I will write the given solution below:

OpenStudy (nikolas):

Using de Moivre's theorem: \[z^{n}+\frac{1}{z^{n}} =(\cos\theta+\imath\sin\theta)^{n}+(\cos\theta+\imath\sin\theta)^{-n}\] \[=\cos n\theta +\imath\sin n\theta +\cos (-n \theta) + \imath\sin(-n \theta)\] \[=\cos n\theta +\imath\sin n\theta +\cos (n \theta) - \imath\sin(n \theta)\] \[=2\cos n\theta\] But I am confused about how to get from the second to the third line. Specifically how does cos(-nθ)+isin(-nθ) become cos(nθ)-isin(nθ)?

OpenStudy (nikolas):

I also need to show that z^n - 1/(z^n) = 2isin(nθ) but hopefully it will be clear once I understand the first problem. Thanks.

OpenStudy (zehanz):

Because cos(-a)=cos(a), for every a, And because sin(-a)=-sin(a) for every a, you can replace them

OpenStudy (zehanz):

(Answering from my mobile phone, typing not easy, srry)

OpenStudy (nikolas):

WOW, thank you so much, not sure why I didn't realize that before but it's so obvious when I think about it. Just goes to show you really need to master the fundamentals before attempting more challenging things. Thanks again.

OpenStudy (zehanz):

Yw!